Question 2. (20 pts.) a) Simplify 2 < (56/))(2 < (-1/3) 2 < (-/ b) Solve z + 2 + 1 - 0

math232
Answer #1

B Z4 + 2 + 1 = 0 (22)+(23)+1 = 0 2 la tatl=0 Where Z²= a Q=-1 -4000 2002 -- I J 1-4 2 -_-1113 2 ܒܝܸܫܵܐ BI 2 - 31 १ 22= -|+ B1